### v1

# SLANG: New Concept Comprehension of Large Language Models

Abstract
The dynamic nature of language, particularly evident in the realm of slang and memes on the Internet, poses serious challenges to the adaptability of large language models (LLMs). Traditionally anchored to static datasets, these models often struggle to keep up with the rapid linguistic evolution characteristic of online communities. This research addresses the critical need to bridge this gap, aiming to enhance LLMs’ comprehension of evolving new concepts on the internet, without the high cost and impracticality of continual retraining. To address this issue, we propose a new benchmark SLANG to assess LLMs’ proficiency in comprehending emerging linguistic trends and a baseline approach FOCUS, which uses causal inference to enhance LLMs to understand new phrases and usage patterns. This approach involves scrutinizing real-world instances of linguistic shifts, serving as contextual beacons, to form more precise and contextually relevant connections between newly emerging expressions and their intended meanings. The empirical analysis shows that our causal inference-based approach outperforms the traditional models in terms of precision and relevance in the interpretation of Internet slang and memes. [^1] Our code is available at https://github.com/Meirtz/FocusOnSlang-Toolbox.
SLANG: New Concept Comprehension of Large Language Models
[…]
1 Introduction
The evolution of human language is the history of human’s innovation and adaptation. From the earliest forms of communication to the complex linguistic systems of today, language has continuously transformed, reflecting, and shaping the human experience. Recently, the language evolution has been accelerated by the developing communication technology, particularly the Internet, which has introduced new dimensions to linguistic evolution (Varis and van Nuenen, 2017; Firth et al., 2019; Hammarström, 2016). These rapid changes in language poses serious challenges to the large language models (LLMs) on understanding the newly emerging concepts (Yang et al., 2023; Sun et al., 2021).
Figure 1: Comparative analysis of LLMs’ understanding of new phrases using CoT Wei et al. (2022) and FOCUS methods. The left side demonstrates the limited understanding through the CoT approach, focusing on the literal interpretation. In contrast, the right side using the FOCUS method shows the model’s enhanced capability to grasp metaphors and deeper meanings.
Generally, LLMs are trained on static data (Brown et al., 2020), which limits their adaptivity to the dynamic and ever-evolving nature of human language. This limitation is particularly pronounced in the context of digital communication, where new forms of expression and knowledge emerge at an unprecedented pace (Sun et al., 2021). The crux of our research lies in exploring the potential of LLMs to adapt to and understand these linguistic transformations and new knowledge without resorting to continual retraining or fine-tuning with updated datasets or retrieving external data.
Note that many popular LLMs are often mainly black-box models. They may make decisions based on superficial patterns rather than justified reasons, which in turn can affect the model’s ability to understand and follow human instructions within context (Tang et al., 2023; Wang et al., 2022, Zhou et al., 2023b, Wang et al., 2023). For example, as shown in Figure 1, while CoT (Chain-of-Thought) Wei et al. (2022) provides a straightforward interpretation of the phrase "The sunset is beautiful, isn’t it?", our method recognizes the underlying metaphorical meaning, identifying it as an acknowledgment of an ending, such as a concluding phase in life or a relationship, in a nuanced conversational exchange. This recognition calls for a more nuanced approach in the development and evaluation of LLMs, which considers not only their performance metrics, but also the underlying principles guiding their decision-making processes.
In this study, we propose FOCUS (Factual cOntext CaUsal AnalysiS), a new methodology centered on causal inference, which can significantly enhance the language model’s capability to understand new phrases. This approach is grounded in the belief that understanding causal relationships in language usage, especially in evolving linguistic contexts, leads to more accurate and adaptable models. Our method involves constructing new datasets from UrbanDictionary Urban Dictionary LLC , comprising both factual and counterfactual instances. These datasets provide diverse linguistic contexts, ranging from common usage to generative scenarios, aiding in the comprehensive understanding of new phrases and concepts.
In addition to FOCUS, we introduce SLANG (Similarity of Lexical Analysis aNd Grasp), a new benchmark designed to evaluate language models’ adaptability to linguistic evolution and vocabulary changes. The SLANG benchmark focuses on assessing the capability of language models to maintain coherence and accuracy in the face of dynamic and unconventional language use, such as slang and idiomatic expressions, thereby evaluating our FOCUS methodology.
2 SLANG
In response to the rapid evolution of the Internet language, we introduce the SLANG benchmark, a tool derived from UrbanDictionary contributions. SLANG aims to evaluate the proficiency of LLMs in interpreting the ever-changing landscape of user-generated content. This benchmark is unique in its division into factual and counterfactual datasets, each serving a distinct purpose in understanding the adaptability of LLMs. The following subsections explore the definition of the SLANG problem, detail the dataset, and discuss the metrics used for evaluation. This structure ensures a comprehensive assessment of LLMs’ capabilities in navigating the complexities of modern linguistic trends.
2.1 Problem Definition
Diving deep into the digital deluge, we encounter the fluid and fleeting world of internet parlance, where slang and idiomatic expressions defy static definitions and demand dynamic interpretation by LLMs. Our journey unfolds with the task of capturing the essence of this linguistic fluidity, mathematically expressed as the challenge of estimating $P(Y|W,X;\mathcal{M})$. This formulation seeks to unravel the likelihood of an explanation $Y$ emerging from a phrase $W$ within its context $X$, as interpreted by the model $\mathcal{M}$.
2.2 Dataset Constructions
Data Extraction
In the construction of the dataset, our methodology involved the systematic extraction of a substantial number of concepts from UrbanDictionary Urban Dictionary LLC , a widely recognized and reliable online repository of Internet slang and colloquial terms. UrbanDictionary stands out for its user-driven content, reflecting current language trends and the evolving nature of the internet lexicon. This platform serves as both a repository and a dynamic community, continuously enriching the evolving landscape of internet language with new concepts. Specifically, our approach involved selecting concepts that were added after a predetermined date, a strategy that was integral to ensuring the relevance and recency of the dataset. The subsequent steps undertaken for the development of this dataset were as follows. Our dataset’s construction commenced with UrbanDictionary as the foundational source. We meticulously extract relevant data, including the phrase, its user-provided definitions, usage examples, and user-generated ratings (upvotes and downvotes). This approach ensured that our dataset was grounded in authentic user-driven language trends. Additionally, the data construction pipeline is set up to automatically include fresh, non-member data for upcoming cutoff dates, ensuring the dataset stays up-to-date and comprehensive.
Figure 2: Comparative histograms illustrating the distribution of upvotes (left) and absolute upvotes (right) across dataset entries. Both histograms are plotted on a logarithmic scale with the vertical axis representing the log percentage of the total dataset and the horizontal axis indicating the log count of upvotes or absolute upvotes. The blue bars represent the raw data, while the orange bars depict the cleaned data, facilitating a direct comparison of the distributions before and after data cleaning.
Content Filtering
The content filtering of our dataset consists of several critical steps, with the foremost being the assurance of content novelty. Millions of concepts on UrbanDictionary have led to a significant proportion of phrases becoming commonplace in everyday communication. This prevalence implies that many of these phrases might already be included in GPT’s training corpus, as seen in the case of widely recognized terms like "ghosting" or "flex", which have transitioned from internet slang to general usage. As such, predicting the meanings of these phrases using LLMs would be redundant. Therefore, our primary objective was to ensure that the phrases and their definitions in our dataset were unfamiliar to the LLMs. In alignment with this objective, we factored in the official knowledge cut-off date of gpt-4-0613, which is April 2021 as detailed in the OpenAI documentation [^2] https://platform.openai.com/docs/models. To ensure the novelty and relevance of our dataset, we selected concepts that were added to UrbanDictionary after January 2022. This temporal gap was strategically chosen to include recent phrases that may have gained popularity online after the cut-off date but were not yet recorded in UrbanDictionary.
An essential aspect of content filtering is to analyze user-generated ratings, which we visualize in Figure 2. Comparative histograms in the figure illustrate the distribution of upvotes (left) and absolute upvotes (right) across dataset entries before and after our cleaning process. Both histograms are plotted on a logarithmic scale with the vertical axis representing the log percentage of the total dataset and the horizontal axis indicating the log count of upvotes or absolute upvotes. The blue bars represent the raw data, while the orange bars depict the cleaned data, facilitating a direct comparison of the distributions before and after data cleaning. This stringent data cleaning was crucial in refining the dataset to ensure the quality and reliability of the entries selected for our research.
We then rigorously employ gpt-4-0613 for an exhaustive filtration process, querying each term repeatedly to ensure that they were not memrozied by the GPT, thereby ensuring the unique and contemporary nature of our dataset.
Additionally, we rigorously filtered inappropriate content to uphold academic integrity, thoroughly examining each entry for explicit, violent, or harmful material to ensure the dataset’s appropriateness and quality.
Factual Dataset
Acknowledging the nonformal nature of the original dataset, we transformed the data into a uniform, formal dictionary format. This transformation involves simplifying the explanations and examples to make them universally understandable while preserving their original online meaning. The transformation process adhered to a custom-designed template, ensuring consistency and clarity across the concepts. However, considering that language models might output synonymous interpretations with varied wording, we augmented each transformed explanation with four additional variants incorporating synonym replacement. This approach was designed to capture the diversity of potential language model outputs. When evaluating predictions, we compare the model’s output with all five versions of the transformed explanations (the original plus four synonym-augmented variants) and select the result that most closely matches. This method ensures a more comprehensive and accurate assessment of the model’s capability to interpret and reproduce the factual content.
Counterfactual Dataset
Another core of our methodology lies in the creation of a counterfactual context. We aim to evaluate the adaptability and interpretative depth of LLMs in semantic decoding and contextual understanding. This approach is characterized by the preservation of original meanings from the factual dataset, while simultaneously introducing entirely new phrases, thereby shifting the linguistic form while maintaining the semantic substance.
This refined strategy requires the innovation of novel phrases that are linguistically distinct from their real-world counterparts, yet congruent with the retained meanings. The challenge lies in crafting these new phrases to be unique in structure, yet resonant with the established meanings, thus presenting a different verbal manifestation of the same conceptual core.
The process involves a creative and analytically rigorous approach, where the rich contextual foundation of the original dataset serves as the semantic basis for the generation of new phrases. These phrases are designed to encapsulate the same ideas as the original meanings, but articulated through a different linguistic prism. Achieving a balance between inventive language use and adherence to logical linguistic norms is crucial to ensuring that new phrases are plausible and coherent within established linguistic frameworks.
Upon generation, these new phrases are systematically paired with their corresponding original meanings, creating a dataset that juxtaposes novel linguistic expressions against established semantic structures. This pairing enables a multifaceted analysis of how LLMs process and comprehend language, specifically assessing their capability to disentangle linguistic form from semantic substance. This methodology offers insight into the nuanced capabilities of LLMs in interpreting language beyond mere recognition, exploring the realms of semantics and contextual linguistics.
2.3 Metrics
Given the nature of this task as a question-answering (QA) problem, traditional QA metrics like F1 score, Recall, and Precision were employed Yang et al. (2018). Moreover, we introduced BLEU (3-gram) Papineni et al. (2002) and ROUGE Lin (2004) as metrics to explore the capacity of the listed methods to generate higher quality interpretations with stricter similarity requirements. Again, considering that language models might output synonymous interpretations with varied wording, we also incorporated sentence similarity measures such as sentence-level similarity and SimCSE Gao et al. (2021), and sentence similarity Team (2023) is obtained by calculating the cosine similarity of the embeddings after the sentences are input into the model all-mpnet-base-v2 (Team, 2023; Song et al., 2020). A positive sample was considered for accuracy calculation if its SimCSE score exceeded 0.7. We also generated five lexically varied yet syntactically and semantically identical interpretations for each dataset entry. To determine the final metric, we selected the interpretation with the highest BLEU score.
3 FOCUS
Figure 3: Structural Causal Model (SCM) of large language models for interpreting new phrases, excluding confounders. The variables $X$ and $W$ encapsulate users’ complex intentions and thoughts, which span intricate emotional expressions, cultural insights, and extensive internet-specific knowledge. Grasping these nuanced aspects directly through an LLM is a challenging endeavor.
This section outlines the novel methodology employed in our research to enhance the adaptability of LLMs in understanding the evolving human language. Our method analogizes the dynamic nature of language to a continuously evolving entity that requires adaptive comprehension strategies.
3.1 Causal Analysis
Structural Causal Models (SCMs) serve as vital tools for elucidating the relationships and influence pathways among variables. We propose a simplified SCM (see Figure 3) to delineate the interpretative processes of an LLM when confronted with novel phrases within their context. IIn this model, users supply both the phrase $W$ and its context $X$, which are then inputted into the LLM, represented by the direct links $X\rightarrow\mathcal{M}\rightarrow Y$ and $W\rightarrow\mathcal{M}\rightarrow Y$, where $\mathcal{M}$ denotes the LLM and $Y$ is the output explanation. These causal links, free from confounders, capture the logical chain from input to interpretation. This SCM sets the stage for our forthcoming discussion where we reintroduce and scrutinize these confounders, thus laying the groundwork for a comprehensive causal analysis.
Figure 4: SCM analysis in FOCUS methodology. This figure presents the complex process of causal inference in the understanding of new phrases, highlighting how the FOCUS approach systematically analyzes and interprets intricate language patterns, emphasizing the causal links between linguistic elements and their interpretive outcomes.
3.1.1 Causal Analysis of Entity
Initially, we reference a typic SCM framework (Wang et al., 2023; Wang et al., 2022), assuming our SCM as $S=\{X,E,Y\}$, where $X$ denotes the context or input, and $E$ represents the confounder, encompassing phrases and other entities within the context. From the perspective of human understanding of new phrases, the interpretation $Y$ is derived from the context $X$ and its entities $E$. The confounder $E$ can be extracted from $X$, leading to the relationships $X\rightarrow Y\leftarrow E$ and $X\rightarrow E$ in our model. By applying the $do$-operation Verma and Pearl (1990), which denoted as $do(X)$, we follow the guidelines (Wang et al., 2023; Wang et al., 2022) to conduct a rigorous causal inference, ensuring that the main effects of the textual context are captured without losing entity information. This operation aims to isolate the effect of the context $X$ on the confounder $E$. Consequently, the relationship between $X$ and $E$ undergoes a change, becoming $\tilde{X}\leftarrow\tilde{E}$, where $\tilde{E}$ represents the modified entity, and $\tilde{X}$ denotes the context obtained after the $do$ operation which effectively substitutes the actual entity.
Figure 5: The four-stage pipeline of FOCUS.
3.1.2 Analysis with Linguistic Factors
In our SCM, denoted as $S=\{X,W,E,Y,R\}$, the output variable $Y$, as an endogenous variable, is influenced by the input context $X$, the input phrase $W$, other entities $E$, and finally, linguistic factors $R$, which include the linguistic structure, style, theme, and cultural background, crucially shape $X$ and constitute the exogenous variables. The model includes direct paths $X\rightarrow Y$ and $W\rightarrow Y$, indicating the immediate influence of context and phrase on the interpretation. The backdoor path $E\leftarrow X\rightarrow W$ and the indirect path $X\rightarrow E\rightarrow W$ demonstrate the mediated effects. The path $R\rightarrow X$ highlights the exogenous influence of linguistic factors on context. This configuration captures the causal mechanisms underlying natural language understanding, emphasizing the direct, indirect, and mediated relationships among the variables. The equation for the causal effect in this context is as follows:
$P(Y=y|X=x)$
$=$ $\sum_{e,r}P(Y=y|X=x,W=w,E=e,R=r)$
$P(W=w|X=x,E=e,R=r)$
$P(E=e)P(R=r)$
In this formula, $P(Y=y|X=x)$ represents the probability of the outcome variable $Y$ being a particular value $y$, given the context $X$ is set to $x$. The summation over $e$ and $r$ encompasses all possible combinations of the values of the entities $E$ and linguistic factors $R$. For each combination, the formula calculates the conditional probability of $Y=y$ given $X=x$, the specific phrase $W=w$, the entity $E=e$, and the linguistic factor $R=r$. This conditional probability is further modulated by the probabilities $P(W=w|X=x,E=e,R=r)$, $P(E=e)$, and $P(R=r)$. These terms represent the likelihood of observing the phrase $W=w$ conditioned on the context, entities, and linguistic factors, as well as the inherent probabilities of the entities $E=e$ and the linguistic factors $R=r$. This comprehensive approach allows for a nuanced understanding of how context, entities, and linguistic factors collectively influence the interpretation $Y$.
Model Method F1 Precision Recall BLUE ROUGE Similarity SimCSE ACC (%)
1 2 L
GPT-3.5 Direct Ouyang et al. (2022) 0.2219 0.1541 0.4320 0.0477 0.2410 0.0547 0.1531 0.1943 0.6806 47.6
CoT Wei et al. (2022) 0.3922 0.3585 0.4583 0.2273 0.4181 0.1822 0.3135 0.5292 0.7653 76.4
CauView Wang et al. (2023) 0.3500 0.2487 0.4885 0.1676 0.3777 0.1506 0.2676 0.5612 0.7868 72.0
FOCUS (Ours) 0.4292 0.4153 0.4524 0.2798 0.4541 0.2131 0.3481 0.5748 0.8017 84.5
GPT-4 Direct Ouyang et al. (2022) 0.2308 0.3474 0.1917 0.0483 0.2597 0.0606 0.1859 0.1616 0.6476 47.2
CoT Wei et al. (2022) 0.4123 0.3947 0.4244 0.2384 0.4370 0.1927 0.3299 0.5521 0.7883 79.3
CauView Wang et al. (2023) 0.3602 0.3444 0.3948 0.1987 0.3917 0.1643 0.3032 0.5515 0.7636 74.7
FOCUS (Ours) 0.4446 0.4280 0.4714 0.3177 0.4721 0.2332 0.3652 0.6032 0.8216 88.2
Table 1: Performance results on the factual dataset.
3.2 Context-based Causal Intervention
Subsequently, we refine our causal intervention approach in the SCM, focusing on the role of context and linguistic factors. To eliminate the entity bias in $W$, it is replaced with a specific placeholder, denoted as $\hat{W}$ (as illustrated in Figure 4 (d)). This alteration aids in isolating the effect of $W$ while excluding entity-specific biases. Further, to concentrate on the contextual content and reduce reliance on shortcuts, $E$ is replaced with $\tilde{E}$, consequently transforming $X$ into $\tilde{X}$. This shift accentuates the role of $R$ in shaping the context, enhancing the model’s capacity to highlight the influence of linguistic factors in a bias-free manner. The updated causal effect formula in the SCM is thus:
$P(Y=y|do(X=x))$
$=$ $\sum_{\tilde{e},r}P(Y=y|X=x,\tilde{W}=\tilde{w},\tilde{E}=\tilde{e},R=r)$
$P(\tilde{W}=\tilde{w}|do(X=x),\tilde{E}=\tilde{e},R=r)$
$P(\tilde{E}=\tilde{e})P(R=r)$
$=$ $\sum_{\tilde{e},r}P(Y=y|X=x,\tilde{W}=\tilde{w},\tilde{E}=\tilde{e},R=r)$
$P(\tilde{W}=\tilde{w}|\tilde{E}=\tilde{e},R=r)P(\tilde{E}=\tilde{e})P(R=r)$
This revised approach and formula ensure a more comprehensive and nuanced analysis of the causal dynamics within the SCM, post-intervention. This approach ensures a more robust and bias-free interpretation within the SCM framework. This equation accounts for the altered relationships in the SCM after $do$-operation. The conditional probabilities and summations are now over the new variables $\tilde{E}$ and $\tilde{W}$, while maintaining the original structure’s intent to adjust for confounding effects and capture the influence of modified entities in the context of $X$.
The core idea of FOCUS is to explore how language models can adhere to guidelines to better understand the content of the context. The goal is to enable LLMs to analyze phrases according to usage examples and provide counterfactual interpretations, thereby understanding the evolving semantics of language. To achieve this objective, we propose a four-stage method, here is a description of each stage and its principles:
Direct Inquiry (DI)
The DI stage is the foundational step in our FOCUS methodology. It primarily involves inputting the usage example (context) $X$ and the phrase $W$ directly into the language model. This stage is represented as $Y_{DI}=\mathcal{M}(X,W,E)$, where $\mathcal{M}$ is the function that defines the processing of the language model. The aim here is to assess the direct effect of the context $X$ on the inferred meaning $Y$ in a straightforward and intervention-free manner. This allows the model to naturally derive the meaning within the given context. Special emphasis is placed on maintaining the integrity of all confounders, ensuring that the phrase’s morphology, literal meaning, and context-based interpretation are coherently analyzed. The DI stage sets the baseline for understanding the meaning of the phrase before any further analytical layers are added in the subsequent stages.
Masked Entity Analysis (MEA)
The MEA stage is a crucial component of our FOCUS methodology, designed to mitigate the impact of the phrase’s literal meaning on its interpretation. In this stage, the phrase $W$ is masked within the context $X$, leading to an evaluation of the meaning $Y$ in the absence of direct influence from $W$. This process is denoted as $Y_{MEA},\hat{W}_{MEA}=\mathcal{M}(X_{\text{masked}},E)$, where $\mathcal{M}$ represents the language model. The primary objective is to understand the meaning based on context, prompting the model to propose synonyms or near-synonyms that could feasibly replace the masked phrase. The MEA stage allows for a nuanced analysis of how the interpretation $Y$ is influenced when $W$ is obscured, yet still impacted by the entities $E$. This stage is instrumental in mitigating the literal interpretation bias of $W$, focusing instead on deriving meaning from the broader linguistic factors present in $X$ and $E$.
Entity Replacement Inquiry (ERI)
The ERI stage plays a pivotal role in our FOCUS methodology, where we explore the impact of varying entity relationships on the interpretation of phrases to delve deeper into context-based interpretation. In this stage, we implement an entity replacement strategy combined with a dropout rate, selectively altering entities within the context $X$. This process, denoted as $Y_{ERI},\hat{W}_{ERI}=\mathcal{M}(\tilde{X}_{\text{masked}},\tilde{E}_{\text{replaced}})$, involves generating multiple context sentences with replaced entities, while a dropout rate determines the extent of entity replacement. The addition of this dropout strategy is not only to introduce randomness but also to prevent the emergence of new entity biases that could arise from completely replacing all entities. By maintaining a mix of original and altered entities, we ensure the robustness of understanding while mitigating the risk of introducing new biases. The language model $\mathcal{M}$ then conducts multiple inferences on these varied contexts to synthesize a cohesive interpretation. This approach enhances the robustness of the model by considering a diverse range of entity interactions. The ERI stage is fundamental in understanding the nuanced effects of entity dynamics within different contexts, providing a more comprehensive causal analysis of phrase interpretation.
Model Method F1 Precision Recall BLUE ROUGE Similarity SimCSE ACC (%)
1 2 L
GPT-3.5 Direct Ouyang et al. (2022) 0.1922 0.1252 0.4497 0.0371 0.2120 0.0477 0.1328 0.1936 0.6051 24.3
CoT Wei et al. (2022) 0.3576 0.3138 0.4407 0.2098 0.3865 0.1713 0.2932 0.4390 0.7232 62.3
CauView Wang et al. (2023) 0.3161 0.2489 0.4471 0.1464 0.3439 0.1297 0.2491 0.4031 0.6954 54.6
FOCUS (Ours) 0.4078 0.4018 0.4294 0.2519 0.4344 0.1955 0.3339 0.5594 0.7836 83.4
GPT-4 Direct Ouyang et al. (2022) 0.2050 0.3018 0.1705 0.0426 0.2341 0.0503 0.1693 0.1645 0.5735 21.0
CoT Wei et al. (2022) 0.3821 0.3573 0.4247 0.2189 0.4091 0.1841 0.3147 0.4383 0.7241 62.6
CauView Wang et al. (2023) 0.3357 0.3024 0.3919 0.1744 0.3701 0.1467 0.2812 0.3815 0.6715 47.2
FOCUS (Ours) 0.4532 0.4598 0.4551 0.3017 0.4763 0.2273 0.3722 0.5645 0.8065 84.9
Table 2: Performance results on the counterfactual dataset.
Final Synthesis (FS)
The FS stage is the integrative phase of the FOCUS methodology, symbolized as $Y_{FS}=\mathcal{M}(Y_{DI},Y_{MEA},Y_{ERI})$. This stage synthesizes insights from Direct Inquiry (DI), Masked Entity Analysis (MEA), and Entity Replacement Inquiry (ERI). It evaluates the interplay between direct, contextually-driven, and variable entity interpretations to forge a comprehensive understanding of the phrase $W$ within context $X$ and its relation to entities $E$. In FS, we assess causal relationships, reconciling the varied interpretations and confounding factors identified in previous stages. This synthesis refines our understanding of language model interpretations, emphasizing contextual richness and linguistic nuance. The FS stage thus concludes the FOCUS methodology, offering a multi-dimensional perspective on language model analytics.
4 Experiments
4.1 Setup
Large Language Models
In our experimental setup, we utilized a subset of 408 samples, filtered from our initial dataset of 7220 concepts, focusing on evaluations over GPT-3.5 (gpt-3.5-turbo-0613) and GPT-4 (gpt-4-0613). These models were selected for their advanced capabilities and the relevance of their knowledge bases, which have a cut-off date prior to April 2021. As we discussed in 2.2, this cut-off is crucial to ensure the novelty and relevancy of our dataset, as it aligns with the temporal frame of our selected UrbanDictionary concepts.
Baselines
Baselines comprised direct inquiry (Direct) Ouyang et al. (2022), Chain-of-Thought (CoT) Wei et al. (2022), CauView Wang et al. (2023), and our FOCUS approaches. In each case, the language model output was parsed and compared with the ground truth. We implement the CauView method in our experiment design by using a two-stage prompt inquiry, due to its lack of a direct inquiry step.
FOCUS
The outputs from the first three stages were processed by a parser that retained only the final conclusions, filtering out the model’s “reasoning" process. In the fourth stage, the language model was briefed about the concise procedures and objectives of the previous three methods. This briefing included potential key information each method might yield and the risks of information omission. Additionally, to control for syntactic variability and minimize error, the prompts for all baselines, including FOCUS, incorporated the same three sample entries for few-shot in-context learning. This ensured that the language model’s output syntax closely resembled that of the dataset.
Experiment F1 Precision Recall BLUE ROUGE-1 ROUGE-2 ROUGE-L Similarity SimCSE ACC (%)
w/o MEA 0.4366 0.4380 0.4484 0.2766 0.4586 0.2059 0.3556 0.8821 0.8014 82.0
w/o ERI 0.4283 0.4300 0.4371 0.2814 0.4593 0.2117 0.3547 0.9021 0.8092 84.0
Table 3: Results of ablation experiments on GPT-4.
4.2 Results
Factual dataset
Our study meticulously evaluated the performance of various language models on the factual dataset, with a particular focus on interpreting new phrases. Table 1 summarizes these findings. We observed that the FOCUS consistently outperformed other approaches across all metrics. Specifically, for the GPT-3.5 model, FOCUS achieved the highest F1 score of 0.4292, Precision of 0.4153, and Recall of 0.4524, along with the best BLUE, ROUGE, Similarity, SimCSE scores, and Accuracy at 84.5%. A similar trend was noted with the GPT-4 model. Here, FOCUS again led the metrics, notably achieving an F1 score of 0.4446, Precision of 0.4280, Recall of 0.4714, and an impressive Accuracy of 88.2%. These results underscore FOCUS’s superior ability to interpret and contextualize slang and idiomatic expressions from UrbanDictionary, as compared to other existing methodologies. The comparative analysis clearly illustrates that our FOCUS approach excels in understanding the nuanced and evolving nature of internet language, thereby marking a significant advancement in the field of natural language processing.
Counterfactual dataset
Following the analysis of the factual dataset, we extended our evaluation to the counterfactual dataset, which presents hypothetical scenarios altering real-world language usage. The performance results, as detailed in Table 2, reveal insightful trends. Notably, our FOCUS method again excelled, outperforming others in metrics with GPT-4, achieving an F1 score of 0.4532, Precision of 0.4598, Recall of 0.4551, and an Accuracy of 84.9%. This reinforces FOCUS’s adeptness at interpreting not only direct language use but also hypothetical, contextually modified scenarios. Comparatively, on the factual dataset, while FOCUS led with substantial margins, the gap in performance was more pronounced in the counterfactual dataset. This discrepancy underscores the enhanced complexity in interpreting altered linguistic contexts, a challenge adeptly met by FOCUS. In essence, our results from both datasets affirm the robustness and versatility of FOCUS in navigating the multifaceted nature of language, effectively bridging the gap between real and hypothetical language use.
4.3 Ablation Study
Our ablation study, focusing on the FOCUS methodology’s components MEA and ERI, utilized GPT-4 on the factual dataset. As shown in Table 3, these components significantly enhance the interpretative capabilities of the model.
MEA (Masked Entity Analysis)
MEA, represented mathematically as $P(Y|W,X,E;\mathcal{M})$, where $Y$ is the interpretation, $W$ the phrase, $X$ the context, and $E$ the entities, critically influences the model’s performance. Its exclusion (w/o MEA) reduced the F1 score to 0.4366 from 0.4446. This result illustrates MEA’s role in disentangling the direct influence of $W$ and $X$ from confounding entities $E$, vital for context-driven interpretation.
ERI (Entity Replacement Inquiry)
The ERI process, which examines the causal links $X\rightarrow Y\leftarrow E$ and $X\rightarrow E$, also shows significant impact. Removing ERI (w/o ERI) decreased the F1 score to 0.4283. ERI’s function in the model, isolating the entity’s influence and exploring alternative causal pathways, proves essential for nuanced language interpretation.
While incorporating either MEA or ERI individually into direct inquiry enhances model performance, their combined use in the FOCUS framework is indispensable for achieving optimal results. This synergy underscores the importance of a comprehensive causal analysis, balancing context and entity dynamics, for the nuanced interpretation of evolving linguistic phenomena.
5 Related Work
5.1 Knowledge Update Methods
LLMs enhance knowledge through Parameter-Efficient Fine-Tuning (PEFT) methods like task-specific parameter addition Houlsby et al. (2019) and low-rank adaptation (LoRA) Hu et al. (2021); Pfeiffer et al. (2020). Despite their efficiency, these methods face challenges such as computational demands, catastrophic forgetting, and reduced task-specific effectiveness Lester et al. (2021). BitFit Ben-Zaken et al. (2021), while simplifying fine-tuning, relies heavily on dataset quality. Trade-offs in computational demands, flexibility, and task compatibility are essential considerations for these methods. Retrieval-Augmented Generation Lewis et al. (2020) and In-context learning-based knowledge editing Zhong et al. (2023) offer dynamic integration of external information, focusing more on fact retrieval than on enhancing deeper understanding.
5.2 Entity Bias and Shortcuts in LLMs
Entity bias Peng et al. (2020); Longpre et al. (2021); Wang et al. (2022, 2023) and shortcuts Du et al. (2021); Saparov and He (2023) in LLMs lead to oversimplified language processing, relying on specific entities or dataset-driven patterns. Entity bias skews model predictions towards certain entities, while shortcuts encompass simplified heuristics, focusing on identifiable features or aspects of the input Du et al. (2022). These patterns limit the models’ understanding and generation of nuanced language, affecting generalization and robustness.
5.3 Causal Intervention Solutions
Causal interventions for debiasing and mitigating shortcuts have gained prominence. Tian et al. (2022) and Zhou et al. (2023a) focus on causal inference and invariant learning for debiasing. CausaLM Feder et al. (2021) provides causal-based model explanations, addressing previous tools’ limitations. Counterfactual methods for debiasing Chen et al. (2023) and eliminating shortcuts Wen et al. (2022) have shown promise, though limitations remain in targeting white-box models and retraining requirements. Our research addresses these limitations through contextual learning and causal inference for bias and shortcut elimination without retraining.
6 Conclusion
In this work, we have explored the dynamic and evolving nature of internet language, particularly slang and memes, and their impact on the adaptability of large language models (LLMs). Our study introduced a novel benchmark, SLANG, to assess LLMs’ proficiency in comprehending emerging linguistic trends. Additionally, we proposed the FOCUS methodology, which utilizes causal inference to enhance understanding of new phrases and usage patterns, going beyond traditional models in precision and relevance. Our analysis revealed that the causal inference-based approach of FOCUS significantly outperforms traditional models in interpreting internet slang and memes. This success is underpinned by our rigorous methodologies involving the construction of datasets from UrbanDictionary and the incorporation of both factual and counterfactual instances to provide diverse linguistic contexts. The results from our experiments demonstrate the enhanced capability of LLMs, equipped with our FOCUS method, to adapt to the rapid evolution of online language. This research marks a significant advancement in the field of natural language processing, highlighting the importance of contextual understanding and adaptability in LLMs for effectively navigating the complexities of evolving human communication.
7 Acknowledge
This paper is partially supported by the National Science Foundation of China under Grant No.U1911401, U21B2046 6237075198.