### v4

# SLANG: New Concept Comprehension of Large Language Models

Abstract
The dynamic nature of language, particularly evident in the realm of slang and memes on the Internet, poses serious challenges to the adaptability of large language models (LLMs). Traditionally anchored to static datasets, these models often struggle to keep up with the rapid linguistic evolution characteristic of online communities. This research aims to bridge this gap by enhancing LLMs’ comprehension of the evolving new concepts on the Internet, without the high cost of continual retraining. In pursuit of this goal, we introduce SLANG, a benchmark designed to autonomously integrate novel data and assess LLMs’ ability to comprehend emerging concepts, alongside FOCUS, an approach uses causal inference to enhance LLMs to understand new phrases and their colloquial context. Our benchmark and approach involves understanding real-world instances of linguistic shifts, serving as contextual beacons, to form more precise and contextually relevant connections between newly emerging expressions and their meanings. The empirical analysis shows that our causal inference-based approach outperforms the baseline methods in terms of precision and relevance in the comprehension of Internet slang and memes. [^1] Our code is available at https://github.com/Meirtz/FocusOnSlang-Toolbox.
SLANG: New Concept Comprehension of Large Language Models
[…]
1 Introduction
Recently, the language evolution has been accelerated by the online community, which has introduced new dimensions to linguistic shifts (Varis and van Nuenen, 2017; Firth et al., 2019; Hammarström, 2016). These rapid changes in language poses serious challenges to the large language models (LLMs) on understanding the newly emerging concepts (Yang et al., 2023; Sun et al., 2021).
Generally, LLMs are trained on static data (Brown et al., 2020), which limits their adaptivity to the dynamic and ever-evolving nature of human language. This limitation is particularly pronounced in the context of digital communication, where new forms of expression and concepts emerge at an unprecedented pace (Sun et al., 2021). Hence, it is essential for LLMs to to understand linguistic shifts and new concepts without constant updates or external data.
Figure 1: Comparative analysis of LLMs’ understanding of new phrases using CoT Wei et al. (2022) and FOCUS methods. The left side demonstrates the limited understanding through the CoT approach, focusing on the literal interpretation. In contrast, the right side using the FOCUS method shows the model’s enhanced capability to grasp metaphors and deeper meanings.
Moreover, LLMs often make decisions based on superficial patterns rather than justified reasons. This can hinder their ability to accurately interpret and follow human instructions, as highlighted in several studies (Tang et al., 2023; Wang et al., 2022, Zhou et al., 2023b, Wang et al., 2023). For example, as depicted in Figure 1, the Chain-of-Thought (CoT) prompting Wei et al. (2022) simply interprets the phrase The sunset is beautiful, isn’t it? and misses the deeper, metaphorical meaning, which could represent an acknowledgment of an ending, like concluding a life phase or a relationship, in a complex conversation. This situation underscores the importance of enhancing and evaluating LLMs in a way that goes beyond their performance metrics. It’s essential to consider the fundamental principles that guide their decision-making processes.
Therefore, we propose SLANG (Similarity of Lexical Analysis aNd Grasp), a benchmark to assess language models’ adaptability to linguistic shifts, and FOCUS (Factual cOntext CaUsal analysiS), an approach based on causal inference for enhancing comprehension of new concepts.
The SLANG benchmark, developed from UrbanDictionary Urban Dictionary LLC , focuses on assessing the capability of language models to maintain coherence and accuracy in the face of dynamic and unconventional language use, such as slang and idiomatic expressions, thereby evaluating the capability of LLMs in grasping new concepts. We select recent entries after a specific cutoff date and filter out phrases already likely in LLM training data. We utilize user-generated ratings (upvotes and downvotes) to refine the dataset, ensuring its quality and comprehensiveness. The dataset is then standardized into a formal dictionary format, simplifying explanations and examples for universal understanding while retaining original meanings. This preprocessing approach ensures SLANG effectively evaluates LLMs’ adaptability to linguistic shifts.
FOCUS employs causal inference to enhance models’ comprehension of new concepts within evolving linguistic contexts. By analyzing causal relationships in language, FOCUS advances models’ predictive capabilities beyond traditional correlation-based learning. This method allows for a nuanced grasp of language dynamics, improving models’ adaptability and effectiveness in applications requiring deep understanding of language use. FOCUS significantly enhanced performance in language model comprehension, demonstrating superior precision and adaptability. With GPT-4, FOCUS achieved an $\text{F}_{1}$ score of 0.4532, precision of 0.4598, and recall of 0.4551, alongside an accuracy of 84.9%, outperforming previous methods in comprehension and adaptability.
The codes for the SLANG and FOCUS toolboxes are open-sourced, contributing to the community’s resources for advancing language model development.
2 SLANG
We introduce SLANG in response to rapidly evolving language. SLANG benchmark evaluates LLMs’ capability to interpret the dynamic landscape of user-generated new concepts. It uniquely features factual and counterfactual datasets, each crucial for gauging LLM adaptability. We detail SLANG’s problem definition, dataset construction, and evaluation metrics in the following subsections.
2.1 Problem
Diving deep into the digital deluge, we encounter the fluid and fleeting world of internet parlance, where slang and idiomatic expressions defy static definitions and demand dynamic interpretation by LLMs. Our journey unfolds with the task of capturing the essence of this linguistic fluidity, mathematically expressed as the challenge of estimating $P(Y|W,X;\mathcal{M})$. This formulation seeks to unravel the likelihood of an explanation $Y$ emerging from a phrase $W$ within its context $X$, as interpreted by the model $\mathcal{M}$.
2.2 Preprocessing
Extraction
Our dataset construction involved extracting numerous concepts from UrbanDictionary, a platform known for user-generated content that reflects current language trends and the evolving internet lexicon, making it a unique, constantly updated repository and dynamic forum for new Internet language concepts. Specifically, our approach involved selecting concepts added after a predetermined date to ensure content novelty. We meticulously extract relevant data, including the phrase, its user-provided definitions, usage examples, and user-generated ratings (upvotes and downvotes). Additionally, the data construction pipeline is set up to automatically include fresh, non-member data for upcoming cutoff dates, ensuring the dataset stays up-to-date and comprehensive.
Filtering
The content filtering of our dataset consists of several steps, with the foremost being the assurance of content novelty. Many UrbanDictionary phrases are now common, potentially pre-included in GPT’s training data. Terms like ghosting or flex illustrate this shift to general usage. To avoid redundancy in LLM predictions, we focused on ensuring that the phrases and definitions in our dataset were new to the LLMs. To solve this, we factored in the official knowledge cut-off date of gpt-4-0613, which is April 2021 as detailed in the OpenAI documentation [^2] https://platform.openai.com/docs/models. To ensure the novelty and relevance of our dataset, we selected concepts that were added to UrbanDictionary after January 2022. This temporal gap was strategically chosen to include recent phrases that may have gained popularity online after the cut-off date but were not yet recorded in UrbanDictionary.
Another essential aspect of content filtering is to analyze user-generated ratings, which we visualize in Figure 2. Comparative histograms in the figure illustrate the distribution of upvotes (left) and absolute upvotes (right) across dataset entries before and after our cleaning process. This stringent data cleaning was crucial in refining the dataset to ensure the quality and reliability of the entries selected for our research.
Additionally, we employed gpt-4-0613 for thorough filtration, ensuring terms were not already memorized by GPT, thus maintaining our dataset’s novelty. We also removed inappropriate content, preserving academic integrity and ensuring quality.
Factual dataset
Acknowledging the informal nature of the original dataset, we transformed the data into a uniform, formal dictionary format. This transformation involves simplifying the explanations and examples to make them universally understandable while preserving their original online meaning. The transformation process adhered to a custom-designed template, ensuring consistency and clarity across the concepts. We also enriched each explanation with four synonym-based variants, ensuring the capture of diverse potential responses. When evaluating predictions, we compare the model’s output with all five versions of the transformed explanations (the original plus four synonym-augmented variants) and select the result that most closely matches. This method ensures a more comprehensive and accurate assessment of the model’s capability to interpret and reproduce the factual content.
Counterfactual dataset
We aim to assess the adaptability and interpretative depth of LLMs in semantic decoding and contextual understanding by preserving original meanings from the factual dataset while introducing new phrases. This approach maintains the semantic substance but shifts the linguistic form. Our strategy involves creating novel phrases that are linguistically distinct yet semantically congruent with their real-world counterparts. Crafting these phrases requires a creative and analytically rigorous approach, leveraging the rich contextual foundation of the original dataset. These new phrases encapsulate the same ideas as the original meanings but are expressed through a different linguistic perspective. Balancing inventive language use with logical linguistic norms is essential to ensure the new phrases are plausible and coherent. Upon generation, these new phrases are paired with their corresponding original meanings, creating a dataset that contrasts novel linguistic expressions against established semantic structures. This pairing allows for a multifaceted analysis of LLMs’ language processing and comprehension, particularly their ability to separate linguistic form from semantic substance. This methodology sheds light on the nuanced capabilities of LLMs in interpreting language beyond mere recognition, delving into semantics and contextual linguistics.
Figure 2: Comparative histograms illustrating the distribution of upvotes (left) and absolute upvotes (right) across dataset entries. Both histograms are plotted on a logarithmic scale with the vertical axis representing the log percentage of the total dataset and the horizontal axis indicating the log count of upvotes or absolute upvotes. The blue bars represent the raw data, while the orange bars depict the cleaned data, facilitating a direct comparison of the distributions before and after data cleaning.
2.3 Metrics
For this task, we employed traditional metrics like $\text{F}_{1}$ score, recall, and precision Yang et al. (2018), and added BLEU (3-gram) Papineni et al. (2002) and ROUGE Lin (2004) for stricter quality checks. Considering that language models might output synonymous interpretations with varied wording, we also incorporated sentence similarity measures such as sentence-level similarity and SimCSE Gao et al. (2021), and sentence similarity Team (2023) as Similarity by calculating the cosine similarity of the embeddings after the sentences are input into the model all-mpnet-base-v2 (Team, 2023; Song et al., 2020). A positive sample was considered for accuracy calculation if its SimCSE score exceeded 0.7. We also generated five lexically varied yet syntactically and semantically identical interpretations for each dataset entry. To determine the final metric, we selected the interpretation with the highest BLEU score.
3 FOCUS
This section outlines the novel approach employed in our research to enhance the adaptability of LLMs in understanding the evolving human language. Our method analogizes the dynamic nature of language to a continuously evolving entity that requires adaptive comprehension strategies.
Figure 3: Structural Causal Model (SCM) of LLMs for interpreting new phrases, excluding confounders. The variables $X$ and $W$ encapsulate users’ complex intentions and thoughts, which span intricate emotional expressions, cultural insights, and extensive internet-specific knowledge. Grasping these nuanced aspects directly through an LLM is a challenging endeavor.
3.1 Causal Analysis
Figure 4: SCM analysis in FOCUS methodology. This figure presents the complex process of causal inference in the understanding of new phrases, highlighting how the FOCUS approach systematically analyzes and interprets intricate language patterns, emphasizing the causal links between linguistic elements and their interpretive outcomes.
Structural Causal Models (SCMs) serve as vital tools for elucidating the relationships and influence pathways among variables. We propose a simplified SCM (see Figure 3) to delineate the interpretative processes of an LLM when confronted with novel phrases within their context. IIn this model, users supply both the phrase $W$ and its context $X$, which are then inputted into the LLM, represented by the direct links $X\rightarrow\mathcal{M}\rightarrow Y$ and $W\rightarrow\mathcal{M}\rightarrow Y$, where $\mathcal{M}$ denotes the LLM and $Y$ is the output explanation. These causal links, free from confounders, capture the logical chain from input to interpretation. This SCM sets the stage for our forthcoming discussion where we reintroduce and scrutinize these confounders, thus laying the groundwork for a comprehensive causal analysis.
Analysis of entity
Initially, we reference a typical SCM framework (Wang et al., 2023; Wang et al., 2022), assuming our SCM as $S=\{X,E,Y\}$, where $X$ denotes the context or input, and $E$ represents the confounder, encompassing phrases and other entities within the context. From the perspective of human understanding of new phrases, the interpretation $Y$ is derived from the context $X$ and its entities $E$. The confounder $E$ can be extracted from $X$, leading to the relationships $X\rightarrow Y\leftarrow E$ and $X\rightarrow E$ in our model. By applying the $do$-operation Verma and Pearl (1990), which denoted as $do(X)$, we follow the guidelines (Wang et al., 2023; Wang et al., 2022) to conduct a rigorous causal inference, ensuring that the main effects of the textual context are captured without losing entity information. This operation aims to isolate the effect of the context $X$ on the confounder $E$. Consequently, the relationship between $X$ and $E$ undergoes a change, becoming $\tilde{X}\leftarrow\tilde{E}$, where $\tilde{E}$ represents the modified entity, and $\tilde{X}$ denotes the context obtained after the $do$ operation which effectively substitutes the actual entity.
Analysis of linguistic factors
In our SCM, denoted as $S=\{X,W,E,Y,R\}$, the output variable $Y$, as an endogenous variable, is influenced by the input context $X$, the input phrase $W$, other entities $E$, and finally, linguistic factors $R$, which include the linguistic structure, style, theme, and cultural background, crucially shape $X$ and constitute the exogenous variables. The model includes direct paths $X\rightarrow Y$ and $W\rightarrow Y$, indicating the immediate influence of context and phrase on the interpretation. The backdoor path $E\leftarrow X\rightarrow W$ and the indirect path $X\rightarrow E\rightarrow W$ demonstrate the mediated effects. The path $R\rightarrow X$ highlights the exogenous influence of linguistic factors on context. This configuration captures the causal mechanisms underlying natural language understanding, emphasizing the direct, indirect, and mediated relationships among the variables. The equation for the causal effect in this context is as follows:
$P(Y=y|X=x)$
$=$ $\sum_{e,r}P(Y=y|X=x,W=w,E=e,R=r)$
$P(W=w|X=x,E=e,R=r)$
$P(E=e)P(R=r)$
In this formula, $P(Y=y|X=x)$ represents the probability of the outcome variable $Y$ being a particular value $y$, given the context $X$ is set to $x$. The summation over $e$ and $r$ encompasses all possible combinations of the values of the entities $E$ and linguistic factors $R$. For each combination, the formula calculates the conditional probability of $Y=y$ given $X=x$, the specific phrase $W=w$, the entity $E=e$, and the linguistic factor $R=r$. This conditional probability is further modulated by the probabilities $P(W=w|X=x,E=e,R=r)$, $P(E=e)$, and $P(R=r)$. These terms represent the likelihood of observing the phrase $W=w$ conditioned on the context, entities, and linguistic factors, as well as the inherent probabilities of the entities $E=e$ and the linguistic factors $R=r$. This comprehensive approach allows for a nuanced understanding of how context, entities, and linguistic factors collectively influence the interpretation $Y$.
3.2 Context-based Causal Intervention
Subsequently, we refine our causal intervention approach in the SCM, focusing on the role of context and linguistic factors. To concentrate on the contextual content and reduce reliance on shortcuts, $E$ is replaced with $\tilde{E}$, consequently transforming $X$ into $\tilde{X}$, (as illustrated in Figure 4 (a) and (b)). Further, to eliminate the entity bias in $W$, it is replaced with a specific placeholder, denoted as $\hat{W}$ (as illustrated in Figure 4 (c) and (d)). This alteration aids in isolating the effect of $W$ while excluding entity-specific biases. Overall, these adjustments accentuate the role of $R$ in shaping the context, enhancing the model’s capacity to highlight the influence of linguistic factors in a bias-free manner. The updated causal effect formula in the SCM is thus:
$P(Y=y|do(X=x))$
$=$ $\sum_{\tilde{e},r}P(Y=y|X=x,\tilde{W}=\tilde{w},\tilde{E}=\tilde{e},R=r)$
$P(\tilde{W}=\tilde{w}|do(X=x),\tilde{E}=\tilde{e},R=r)$
$P(\tilde{E}=\tilde{e})P(R=r)$
$=$ $\sum_{\tilde{e},r}P(Y=y|X=x,\tilde{W}=\tilde{w},\tilde{E}=\tilde{e},R=r)$
$P(\tilde{W}=\tilde{w}|\tilde{E}=\tilde{e},R=r)P(\tilde{E}=\tilde{e})P(R=r)$
This revised approach and formula ensure a more comprehensive and nuanced analysis of the causal dynamics within the SCM, post-intervention. This approach ensures a more robust and bias-free interpretation within the SCM framework. This equation accounts for the altered relationships in the SCM after $do$-operation. The conditional probabilities and summations are now over the new variables $\tilde{E}$ and $\tilde{W}$, while maintaining the original structure’s intent to adjust for confounding effects and capture the influence of modified entities in the context of $X$.
Figure 5: The four-stage pipeline of FOCUS.
The core idea of FOCUS is to explore how language models can adhere to guidelines to better understand the content of the context. The goal is to enable LLMs to analyze phrases according to usage examples and provide counterfactual interpretations, thereby understanding the evolving semantics of language. To achieve this objective, we propose a four-stage method (as shown in Figure 5), each with its principles:
Direct Inquiry (DI)
In the Direct Inquiry (DI) stage, we input the usage example (context) $X$ and phrase $W$ into the language model. Represented as $Y_{DI}=\mathcal{M}(X,W,E)$, this stage aims to evaluate the direct effect of $X$ on the inferred meaning $Y$. It allows the model to derive meaning naturally within the given context, focusing on the coherence of the phrase’s morphology, literal meaning, and context-based interpretation. DI sets a baseline for understanding the phrase’s meaning before adding more analytical layers.
Masked Entity Inquiry (MEI)
In the Masked Entity Inquiry (MEI) stage, we mask the phrase $W$ within context $X$ to analyze the meaning $Y$ without $W$’s direct influence. This process, $Y_{MEA},\hat{W}{MEA}=\mathcal{M}(X{\text{masked}},E)$, helps the model suggest synonyms or near-synonyms for the masked phrase. MEI focuses on extracting meaning from broader linguistic factors in $X$ and $E$, reducing bias towards $W$’s literal interpretation and enhancing context-based understanding.
Model Prompting Method $\textbf{F}_{\textbf{1}}$ Precision Recall BLUE ROUGE Similarity SimCSE ACC (%)
1 2 L
GPT-3.5 Direct Ouyang et al. (2022) 0.2219 0.1541 0.4320 0.0477 0.2410 0.0547 0.1531 0.1943 0.6806 47.6
CoT Wei et al. (2022) 0.3922 0.3585 0.4583 0.2273 0.4181 0.1822 0.3135 0.5292 0.7653 76.4
CauView Wang et al. (2023) 0.3500 0.2487 0.4885 0.1676 0.3777 0.1506 0.2676 0.5612 0.7868 72.0
FOCUS (Ours) 0.4292 0.4153 0.4524 0.2798 0.4541 0.2131 0.3481 0.5748 0.8017 84.5
GPT-4 Direct Ouyang et al. (2022) 0.2308 0.3474 0.1917 0.0483 0.2597 0.0606 0.1859 0.1616 0.6476 47.2
CoT Wei et al. (2022) 0.4123 0.3947 0.4244 0.2384 0.4370 0.1927 0.3299 0.5521 0.7883 79.3
CauView Wang et al. (2023) 0.3602 0.3444 0.3948 0.1987 0.3917 0.1643 0.3032 0.5515 0.7636 74.7
FOCUS (Ours) 0.4446 0.4280 0.4714 0.3177 0.4721 0.2332 0.3652 0.6032 0.8216 88.2
Table 1: Performance results on the factual dataset.
Entity Replacement Inquiry (ERI)
In the Entity Replacement Inquiry (ERI) stage, we alter entities in context $X$ to assess phrase interpretation variability. We use a dropout rate for entity alteration in $Y_{ERI},\hat{W}_{ERI}=\mathcal{M}(\tilde{X}_{\text{masked}},\tilde{E}_{\text{replaced}})$. This introduces a balance of original and new entities, enhancing model robustness without bias. ERI helps understand entity dynamics’ effects on interpretation, providing deeper causal analysis.
Synthesis (SY)
In the Synthesis (SY) stage, we integrate insights from Direct Inquiry, Masked Entity Inquiry, and Entity Replacement Inquiry. Represented as $Y_{FS}=\mathcal{M}(Y_{DI},Y_{MEA},Y_{ERI})$, this phase evaluates the interplay between direct, contextual, and entity-variable interpretations. SY reconciles varied interpretations and confounders, refining the model’s understanding of language nuances. This final stage offers a multi-dimensional perspective on language model analytics, emphasizing contextual richness.
4 Experiments
4.1 Setup
Large Language Models
In our experimental setup, we utilized a subset of 408 samples, filtered from our initial dataset of 7220 concepts, focusing on evaluations over GPT-3.5 (gpt-3.5-turbo-0613) and GPT-4 (gpt-4-0613). These models were selected for their advanced capabilities and the relevance of their knowledge bases, which have a cut-off date prior to April 2021. As we discussed in 2.2, this cut-off is crucial to ensure the novelty and relevancy of our dataset, as it aligns with the temporal frame of our selected UrbanDictionary concepts.
Baselines
Baselines comprised direct inquiry (Direct) Ouyang et al. (2022), Chain-of-Thought (CoT) Wei et al. (2022), CauView Wang et al. (2023), and our FOCUS approaches. In each case, the language model output was parsed and compared with the ground truth. We implement the CauView method in our experiment design by using a two-stage prompt inquiry, due to its lack of a direct inquiry step.
4.2 Results
Results on factual dataset
Our study meticulously evaluated the performance of various language models on the factual dataset, with a particular focus on interpreting new phrases. Table 1 summarizes these findings. We can see that the FOCUS consistently outperforms other approaches across all metrics. Specifically, for the GPT-3.5 model, FOCUS achieves the highest $\text{F}_{1}$ score of 0.4292, precision of 0.4153, and recall of 0.4524, along with the best BLUE, ROUGE, similarity, SimCSE scores, and accuracy at 84.5%. A similar trend is noted with the GPT-4 model. Here, FOCUS again lead the metrics, notably achieving an $\text{F}_{1}$ score of 0.4446, precision of 0.4280, recall of 0.4714, and an impressive accuracy of 88.2%. These results underscore FOCUS’s superior ability to interpret and contextualize slang and idiomatic expressions from UrbanDictionary, as compared to other existing methodologies.
Model Prompting Method $\textbf{F}_{\textbf{1}}$ Precision Recall BLUE ROUGE Similarity SimCSE ACC (%)
1 2 L
GPT-3.5 Direct Ouyang et al. (2022) 0.1922 0.1252 0.4497 0.0371 0.2120 0.0477 0.1328 0.1936 0.6051 24.3
CoT Wei et al. (2022) 0.3576 0.3138 0.4407 0.2098 0.3865 0.1713 0.2932 0.4390 0.7232 62.3
CauView Wang et al. (2023) 0.3161 0.2489 0.4471 0.1464 0.3439 0.1297 0.2491 0.4031 0.6954 54.6
FOCUS (Ours) 0.4078 0.4018 0.4294 0.2519 0.4344 0.1955 0.3339 0.5594 0.7836 83.4
GPT-4 Direct Ouyang et al. (2022) 0.2050 0.3018 0.1705 0.0426 0.2341 0.0503 0.1693 0.1645 0.5735 21.0
CoT Wei et al. (2022) 0.3821 0.3573 0.4247 0.2189 0.4091 0.1841 0.3147 0.4383 0.7241 62.6
CauView Wang et al. (2023) 0.3357 0.3024 0.3919 0.1744 0.3701 0.1467 0.2812 0.3815 0.6715 47.2
FOCUS (Ours) 0.4532 0.4598 0.4551 0.3017 0.4763 0.2273 0.3722 0.5645 0.8065 84.9
Table 2: Performance results on the counterfactual dataset.
Experiment $\textbf{F}_{\textbf{1}}$ Precision Recall BLUE ROUGE-1 ROUGE-2 ROUGE-L Similarity SimCSE ACC (%)
w/o MEA 0.4366 0.4380 0.4484 0.2766 0.4586 0.2059 0.3556 0.8821 0.8014 82.0
w/o ERI 0.4283 0.4300 0.4371 0.2814 0.4593 0.2117 0.3547 0.9021 0.8092 84.0
Table 3: Results of ablation experiments on GPT-4.
Results on counterfactual dataset
Following the analysis of the factual dataset, we extended our evaluation to the counterfactual dataset, which presents hypothetical scenarios altering real-world language usage. The performance results, as detailed in Table 2, reveal insightful trends. Notably, our FOCUS method again outperforms others in metrics with GPT-4, achieving an $\text{F}_{1}$ score of 0.4532, precision of 0.4598, recall of 0.4551, and an accuracy of 84.9%. This reinforces FOCUS’s adeptness at interpreting not only direct language use but also hypothetical, contextually modified scenarios. In essence, our results from both datasets affirm the robustness and versatility of FOCUS in navigating the multifaceted nature of language, effectively bridging the gap between real and hypothetical language use.
4.3 Ablation Study
Our ablation study, focusing on the FOCUS methodology’s components MEI and ERI, utilized GPT-4 on the factual dataset. As shown in Table 3, these components significantly enhance the interpretative capabilities of the model.
MEI
The MEI stage, represented mathematically as $P(Y|W,X,E;\mathcal{M})$, where $Y$ is the interpretation, $W$ the phrase, $X$ the context, and $E$ the entities, critically influences the model’s performance. Its exclusion (w/o MEI) reduced the $\text{F}_{1}$ score to 0.4366 from 0.4446. This result illustrates MEA’s role in disentangling the direct influence of $W$ and $X$ from confounding entities $E$, vital for context-driven interpretation.
ERI
The ERI stage, which examines the causal links $X\rightarrow Y\leftarrow E$ and $X\rightarrow E$, also shows significant impact. Removing ERI (w/o ERI) decreased the $\text{F}_{1}$ score to 0.4283. ERI’s function in the model, isolating the entity’s influence and exploring alternative causal pathways, proves essential for nuanced language interpretation.
While incorporating either MEI or ERI individually into direct inquiry enhances model performance, their combined use in the FOCUS framework is indispensable for achieving optimal results. This synergy underscores the importance of a comprehensive causal analysis, balancing context and entity dynamics, for the nuanced interpretation of evolving linguistic phenomena.
5 Related Work
5.1 Knowledge Update Methods
LLMs enhance knowledge through parameter-efficient fine-tuning methods like task-specific parameter addition Houlsby et al. (2019) and low-rank adaptation (LoRA) Hu et al. (2021); Pfeiffer et al. (2020). Despite their efficiency, these methods face challenges such as computational demands, catastrophic forgetting, and reduced task-specific effectiveness Lester et al. (2021). BitFit Ben-Zaken et al. (2021), while simplifying fine-tuning, relies heavily on dataset quality. Trade-offs in computational demands, flexibility, and task compatibility are essential considerations for these methods. retrieval-augmented generation Lewis et al. (2020) and in-context learning-based knowledge editing Zhong et al. (2023) offer dynamic integration of external information, focusing more on fact retrieval than on enhancing deeper understanding.
5.2 Entity Bias and Shortcuts in LLMs
Entity bias Peng et al. (2020); Longpre et al. (2021); Wang et al. (2022, 2023) and shortcuts Du et al. (2021); Saparov and He (2023) in LLMs lead to oversimplified language processing, relying on specific entities or dataset-driven patterns. Entity bias skews model predictions towards certain entities, while shortcuts encompass simplified heuristics, focusing on identifiable features or aspects of the input Du et al. (2022). These patterns limit the models’ understanding and generation of nuanced language, affecting generalization and robustness.
5.3 Causal Intervention Solutions
Causal interventions for debiasing and mitigating shortcuts have gained prominence. Tian et al. (2022) and Zhou et al. (2023a) focus on causal inference and invariant learning for debiasing. CausaLM Feder et al. (2021) provides causal-based model explanations, addressing previous tools’ limitations. Counterfactual methods for debiasing Chen et al. (2023) and eliminating shortcuts Wen et al. (2022) have shown promise, though limitations remain in targeting white-box models and retraining requirements. Our research addresses these limitations through contextual learning and causal inference for bias and shortcut elimination without retraining.
6 Conclusion
In this work, we have explored the dynamic and evolving nature of internet language, particularly slang and memes, and their impact on the adaptability of large language models (LLMs). Our study introduced a novel benchmark, SLANG, to assess LLMs’ proficiency in comprehending emerging linguistic trends. Additionally, we proposed the FOCUS methodology, which utilizes causal inference to enhance understanding of new concepts, going beyond other methods in precision and relevance. This success is underpinned by our rigorous methodologies involving the construction of datasets from UrbanDictionary and the incorporation of both factual and counterfactual instances to provide diverse linguistic contexts. The results from our experiments demonstrate the enhanced capability of LLMs, equipped with our FOCUS method, to adapt to the rapid evolution of online language. This research marks a significant advancement in the field of natural language processing, highlighting the importance of contextual understanding and adaptability in LLMs for effectively navigating the complexities of evolving human communication.
Limitations
The main limitation of this work lies in its reliance on Internet-specific data, primarily sourced from UrbanDictionary. While this provides a rich and current linguistic landscape, it may not comprehensively represent the vast diversity of global Internet slang and memes, potentially limiting the generalizability of our findings. Additionally, the performance of FOCUS heavily depends on the quality and variety of the dataset used. Our approach, although innovative, does not fully address the complexities of linguistic evolution in non-English or morphologically rich languages. Future work should explore more diverse datasets and extend our methodology to other languages and linguistic contexts.
Ethics Statement
Our research acknowledges that while methods like the SLANG benchmark and FOCUS approach enhance LLM’s understanding of Internet language, they cannot entirely eliminate the propagation of harmful content. Users must exercise caution and cultural sensitivity, especially when interpreting slang and memes, to avoid reinforcing stereotypes or biases. Our work encourages responsible use, emphasizing the importance of respecting diverse linguistic origins and the natural evolution of language.
Acknowledge
This paper is partially supported by the National Science Foundation of China under Grant No.U1911401, U21B2046 6237075198.